The shelf life, in days, for bottles of a certain prescribed medicine is a random variable having the density function below. Find the probability that a bottle of this medicine will have a shell life of

f(x)= {20000/(x+100)^3, x>0,
.........0,(included on the bracket) elsewhere.

(a) at least 365 days. (Decimal form)
(b) at most 365 days. (Decimal form)
1
Expert's answer
2020-12-10T13:04:31-0500

a. Let X be the shelf life for a bottle of the medicine, i.e the random variable with the density function f.

"P(x\geq365)""\space =\int^{\infin}_{x=365}f(x)"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=\int^{\infin}_{x=365}\frac{20000}{(x+100)^3}"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=(-\frac{10000}{(x+100)^2})\mid^{\infin}_{365}"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=0.04624812117"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=0.0462"

b.


"P(x\leq365)=1-P(x\geq365)"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=1-\int^{\infin}_{x=365}f(x)"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=1-\int^{\infin}_{x=365}\frac{20000}{(x+100)^3}"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=1-(-\frac{10000}{(x+100)^2})\mid^{\infin}_{365}"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=1-\frac{10000}{(365+100)^2}"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=1-\frac{400}{8649}"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=0.95375187883"


"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=0.9538"